# Feature learning in finite-width Bayesian deep linear networks with multiple outputs and convolutional layers

### v1

@@ A.2 Proofs for the FC-DLN
$\mathbf{S}_{1:P}|U_{1:L}^{\top}U_{1:L}\sim\mathcal{N}(0,U_{1:L}^{\top}U_{1:L}\otimes(\lambda^{*}N_{0})^{-1}X^{\top}X).$
and the thesis follows. $\blacklozenge$
@@
Let $\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})$ be as in Eq. (8) and $Q^{(L)}=Q^{(L)}(Q_{1},\dots,Q_{L})$ be defined as in Proposition 1. Let $p_{\mathrm{post}}(\mathbf{s}_{0},\dots,\mathbf{s}_{P}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)$ be the posterior distribution of $(\mathbf{S}_{0},\dots,\mathbf{S}_{P})$ given $\mathbf{y}_{1:P}$ in a FC-DLN. If $\det(\tilde{X}^{\top}\tilde{X})>0$, then
$p_{\mathrm{post}}(\mathbf{s}_{0},\mathbf{s}_{1},\dots,\mathbf{s}_{P}|\mathbf{y}_{1:P},\tilde{X})=\int_{(\mathcal{S}^{+}_{D})^{L}}\frac{e^{-\frac{1}{2}(\mathbf{s}-\mathbf{m})^{\top}(\beta\Pi_{0}+\Sigma^{-1})(\mathbf{s}-\mathbf{m})}}{(2\pi)^{\frac{D(P+1)}{2}}\det((\beta\Pi_{0}+\Sigma^{-1})^{-1})^{\frac{1}{2}}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\tilde{X},\beta)$
@@
$\Sigma=\begin{pmatrix}\Sigma_{00}&\Sigma_{01}\\ \Sigma_{01}^{\top}&\Sigma_{11}\\ \end{pmatrix}:=Q^{(L)}\otimes\frac{\tilde{X}^{\top}\tilde{X}}{N_{0}\lambda^{*}},\qquad\Pi_{0}=\begin{pmatrix}0_{D\times D}&0_{DP\times P}\\ 0_{P\times DP}&\mathbb{1}_{DP}\\ \end{pmatrix},\quad\mathbf{m}:=\begin{pmatrix}\mathbf{m}_{0}\\ \mathbf{m}_{1}\\ \end{pmatrix}=\begin{pmatrix}\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\\ \Sigma_{11}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\\ \end{pmatrix}\mathbf{y}_{1:P},$
@@
with $\Sigma:=Q^{(L)}\otimes(N_{0}\lambda^{*})^{-1}\tilde{X}^{\top}\tilde{X}$. Since $\tilde{X}^{\top}\tilde{X}$ is assumed to be strictly positive definite and hence invertible, using that $(A\otimes B)^{-1}=A^{-1}\otimes B^{-1}$ (see, e.g., Proposition 1.28 [55]) it follows that also $\Sigma$ is invertible. Simple computations show that
$\begin{split}{\beta}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})^{\top}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})+\mathbf{s}\Sigma^{-1}\mathbf{s}&=(\mathbf{s}-\mathbf{y}^{0})^{\top}{\Pi_{0}}{\beta}(\mathbf{s}-\mathbf{y}^{0})+\mathbf{s}\Sigma^{-1}\mathbf{s}\\ &=(\mathbf{s}-\mathbf{m})^{\top}({\Pi_{0}}{\beta}+\Sigma^{-1})(\mathbf{s}-\mathbf{m})+\varphi(\Sigma,\mathbf{y}_{1:P})\\ \end{split}$
@@
$\begin{split}&\mathbf{y}^{0}:=(0_{K}^{\top},\mathbf{y}_{1:P}^{\top})^{\top},\quad\mathbf{m}=({\Pi_{0}}{\beta}+\Sigma^{-1})^{-1}{\Pi_{0}}{\beta}\mathbf{y}^{0}={\beta}({\Pi_{0}}{\beta}+\Sigma^{-1})^{-1}\mathbf{y}^{0}\,,\\ &\varphi(\Sigma,\mathbf{y}_{1:P})=\beta{\mathbf{y}_{1:P}^{\top}\mathbf{y}_{1:P}}-(\mathbf{m})^{\top}({\Pi_{0}}{\beta}+\Sigma^{-1})\mathbf{m}\,.\\ \\ \end{split}$
@@
To conclude it remains to compute the previous matrix. We have already noted that, by $\tilde{X}^{\top}\tilde{X}>0$, it follows that also $\Sigma$ is invertible. Being $X^{\top}X$ and $\mathbf{x}_{0}^{\top}\mathbf{x}_{0}$ strictly positive, then the Schur complement of $\Sigma_{00}$ in $\Sigma$, $\Sigma^{*}_{00}:=\Sigma_{00}-\Sigma_{01}\Sigma_{11}^{-1}\Sigma_{01}^{\top}$, is well-defined and strictly positive (Proposition 1.34 [55]). In particular, using the Banachiewicz inversion formula, one can write the inverse of $\Sigma$ as
@@
See, e.g., Theorem 1.2 in [65]. Using again the Banachiewicz inversion formula on $\beta\Pi_{0}+\Sigma^{-1}$ (which is again strictly positive), after some tedious computations, one gets that
@@
$\Sigma_{00}=Q^{(L)}\otimes\mathbf{x}_{0}^{\top}\mathbf{x}_{0},\quad\Sigma_{01}=Q^{(L)}\otimes\mathbf{x}_{0}^{\top}X,\quad\Sigma_{11}=Q^{(L)}\otimes X^{\top}X,$
@@
Proof of Proposition 5. We give the proof in the case $L=1$, the general case follows the same lines, or using Exercise 4.2.7 in [59] and the independence of the components. The result is a consequence of a very general version of the Cramér’s theorem. See Theorem 6.1.3 and Corollary 6.1.6 in [59]. Recall that if $Q_{1}\sim\mathcal{W}_{D}(\cdot|\mathbb{1}_{D},N)$ then $Q_{\ell}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}$ where $\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}\sim\mathcal{W}_{D}(\cdot|\mathbb{1}_{D},1)$ are independent and identically distributed taking values in $\mathcal{S}^{+}_{D}$. Now $\mathcal{S}^{+}_{D}$ is a convex open cone in the (topological) vector space $\mathcal{S}_{D}$ of the $D\times D$ symmetric matrices, endowed with the scalar product $(A,B)=\operatorname{tr}(AB)$. The space $\mathcal{S}_{D}$ is locally convex, Hausdorff and the closure of $\mathcal{S}^{+}_{D}$ in $\mathcal{S}_{D}$ is convex and separable and complete (with the topology induced by $\mathcal{S}_{D})$. Clearly $P\{\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}\in\overline{\mathcal{S}^{+}_{D}}\}=1$ and any closed convex hull of a compact in $\overline{\mathcal{S}^{+}_{D}}$ is compact. Hence all the hypotheses of the General Cramér’s theorem given in Chapter 6 of [59] are satisfied and hence $Q_{\ell}$ satisfies a (weak) LDP (when $N\to+\infty$) with rate function given by $I(Q)=\Lambda^{*}(Q):=\sup_{A\in\mathcal{S}_{D}}\{\operatorname{tr}(QA)-\Lambda(A)\}$ where
@@
In point of fact, see Corollary 6.1.6 in [59], it holds also a strong LDP provided that 00 is in the interior of the domain of $\Lambda$, as we shall show. The well-known expression for the Laplace transform of a Wishart distribution with one degree of freedom (see (20)) gives $\Lambda(A)=-\frac{1}{2}\log(\det(\mathbb{1}_{D}-2A))$ wherever the eigenvalues of $\mathbb{1}_{D}-2A$ are positive and +$\infty$+ ∞ otherwise. In computing $\Lambda^{*}(Q)$ we can restrict $A$ to be such that the eigenvalues of $\mathbb{1}_{D}-2A$ are positive, since otherwise $-\Lambda(A)=-\infty$. Now we check that
@@
Proof. Proposition 6    The proof follows by Lemma 8. To check the assumptions, note that
$Q_{1},\dots,Q_{L}\mapsto\Sigma(Q_{1},\dots,Q_{L})=\begin{pmatrix}\Sigma_{00}&\Sigma_{01}\\ \Sigma_{01}^{\top}&\Sigma_{11}\\ \end{pmatrix}:=Q^{(L)}\otimes\frac{\tilde{X}^{\top}\tilde{X}}{N_{0}\lambda^{*}}$
@@
hence $\Sigma_{11}\mapsto\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$ is bounded on bounded set of $\Sigma_{11}$. Using the continuity of $(Q_{1},\dots,Q_{L})\mapsto\Sigma_{11}(Q_{1},\dots,Q_{L})$ it follows that also $(Q_{1},\dots,Q_{L})\mapsto\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$ is bounded when $\sum_{\ell=1}^{L}\|Q_{\ell}\|_{2}\leq K$. It remains to check that
@@
Letting $\lambda_{i\ell}$ the eigenvalues of $Q_{\ell}$, the claim follows since $\operatorname{tr}(Q_{\ell})-\log(\det(Q_{\ell}))=\sum_{i=1}^{P}\Big{(}\lambda_{i}+\log(\lambda_{i})\Big{)}$. $\blacklozenge$
@@ A.3 Proofs for the C-DLN
$\bar{\mathbf{s}}_{1:P}^{\top}\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})\bar{\mathbf{s}}_{1:P}=\sum_{\mu,\nu}\sum_{a_{0}}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},s}}{\lambda^{*}C_{0}}\bar{s}_{\mu}\bar{s}_{\nu}=\sum_{a_{0}}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}A^{*}_{r,a_{0}}A^{*}_{s,a_{0}}$
@@
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ $=\mathbf{E}\Biggl{[}\det\left((\mathbb{1}_{C_{L}}\otimes K^{(L-1)})^{-1}+\frac{1}{\lambda_{L}C_{L}N_{0}}\mathbb{1}_{C_{L}}\otimes\mathbb{1}_{N_{0}}\otimes(\bar{\mathbf{s}}_{1:P}\bar{\mathbf{s}}_{1:P}^{\top})\right)^{-\frac{1}{2}}\det(\mathbb{1}_{C_{L}}\otimes K^{(L-1)})^{-\frac{1}{2}}\Biggr{]}$ (32)
@@
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ $=\mathbf{E}\int_{\mathcal{S}^{+}_{N_{0}}}\exp\Biggl{[}-\frac{1}{2\lambda_{L}N_{0}}\sum_{i,j}Q_{L,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}K^{(L-1)}_{ij,\mu\nu}\Biggr{]}\mathcal{W}_{N_{0}}\Big{(}dQ_{L}|\frac{1}{C_{L}}\mathbb{1}_{N_{0}},C_{L}\Big{)}$ (33)
$=\mathbf{E}\int_{\mathcal{S}^{+}_{N_{0}}}\exp\Biggl{[}-\frac{1}{2\lambda_{L}\lambda_{L-1}C_{L-1}N_{0}}\sum_{i,j}Q^{*}_{L,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}\sum_{a_{L-1}=1}^{C_{L-1}}h^{(L-1)}_{a_{L-1},i,\mu}h^{(L-1)}_{a_{L-1},j,\nu}\Biggr{]}\mathcal{W}_{N_{0}}\Big{(}dQ_{L}|\frac{1}{C_{L}}\mathbb{1}_{N_{0}},C_{L}\Big{)}$
@@
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ $=\mathbf{E}\int_{(\mathcal{S}^{+}_{N_{0}})^{2}}\exp\Biggl{[}-\frac{1}{2\lambda_{L}\lambda_{L-1}N_{0}}\operatorname{tr}\Biggl{(}Q_{L-1}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}[U^{*}_{L}K^{(L-2)}_{\cdot,\mu\nu}(U^{*}_{L})^{\top}]\Biggr{)}\Biggl{]}\prod_{\ell\in\{L,L-1\}}\mathcal{W}_{N_{0}}(dQ_{\ell}|\frac{1}{C_{\ell}}\mathbb{1}_{N_{0}},C_{\ell})$ (36)
$=\mathbf{E}\int_{(\mathcal{S}^{+}_{N_{0}})^{2}}\exp\Biggl{[}-\frac{1}{2\lambda_{L}\lambda_{L-1}\lambda_{L-2}C_{L-2}N_{0}}\sum_{i,j}Q^{*}_{L-1,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}\sum_{a_{L-2}=1}^{C_{L-2}}h^{(L-2)}_{a_{L-2},i,\mu}h^{(L-2)}_{a_{L-2},j,\nu}\Biggr{]}$
×$\hskip 300.00046pt\times\prod_{\ell\in\{L,L-1\}}\mathcal{W}_{N_{0}}(dQ_{\ell}|\frac{1}{C_{\ell}}\mathbb{1}_{N_{0}},C_{\ell})\,,$

### v2

@@ A.2 Proofs for the FC-DLN
$\mathbf{S}_{1:P}|U_{1:L}^{\top}U_{1:L}\sim\mathcal{N}(0,(\lambda^{*}N_{0})^{-1}X^{\top}X\otimes U_{1:L}^{\top}U_{1:L}).$
and the thesis follows.
$\blacklozenge$
@@
Let $\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})$ be as in Eq. (8) and $Q^{(L)}=Q^{(L)}(Q_{1},\dots,Q_{L})$ be defined as in Proposition 1. Let $p_{\mathrm{post}}(\mathbf{s}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)$ be the posterior distribution of $(\mathbf{S}_{0},\dots,\mathbf{S}_{P})$ given $\mathbf{y}_{1:P}$ in a FC-DLN. If $\det(\tilde{X}^{\top}\tilde{X})>0$, then
$p_{\mathrm{post}}(\mathbf{s}|\mathbf{y}_{1:P},\tilde{X})\\ =\int_{(\mathcal{S}^{+}_{D})^{L}}\frac{e^{-\frac{1}{2}(\mathbf{s}-\mathbf{m})^{\top}(\beta\Pi_{0}+\Sigma^{-1})(\mathbf{s}-\mathbf{m})}}{(2\pi)^{\frac{D(P+1)}{2}}\det((\beta\Pi_{0}+\Sigma^{-1})^{-1})^{\frac{1}{2}}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\tilde{X},\beta)$
@@
$\Sigma=\begin{pmatrix}\Sigma_{00}&\Sigma_{01}\\ \Sigma_{01}^{\top}&\Sigma_{11}\\ \end{pmatrix}:=\frac{\tilde{X}^{\top}\tilde{X}}{N_{0}\lambda^{*}}\otimes Q^{(L)},\qquad\Pi_{0}=\begin{pmatrix}0_{D\times D}&0_{P\times DP}\\ 0_{DP\times P}&\mathbb{1}_{DP}\\ \end{pmatrix}$
$\mathbf{m}:=\begin{pmatrix}\mathbf{m}_{0}\\ \mathbf{m}_{1}\\ \end{pmatrix}=\begin{pmatrix}\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\\ \Sigma_{11}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\\ \end{pmatrix}\mathbf{y}_{1:P},$
@@
with $\Sigma:=(N_{0}\lambda^{*})^{-1}\tilde{X}^{\top}\tilde{X}\otimes Q^{(L)}$. Since $\tilde{X}^{\top}\tilde{X}$ is assumed to be strictly positive definite and hence invertible, using that $(A\otimes B)^{-1}=A^{-1}\otimes B^{-1}$ (see, e.g., Proposition 1.28 Eaton, 2007) it follows that also $\Sigma$ is invertible. Simple computations show that
$\begin{split}{\beta}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})^{\top}(\mathbf{s}_{1:P}-\mathbf{y}_{1:P})+\mathbf{s}\Sigma^{-1}\mathbf{s}&=(\mathbf{s}-\tilde{\mathbf{y}})^{\top}{\Pi_{0}}{\beta}(\mathbf{s}-\tilde{\mathbf{y}})+\mathbf{s}\Sigma^{-1}\mathbf{s}\\ &=(\mathbf{s}-\mathbf{m})^{\top}({\Pi_{0}}{\beta}+\Sigma^{-1})(\mathbf{s}-\mathbf{m})+\varphi(\Sigma,\mathbf{y}_{1:P})\\ \end{split}$
@@
$\begin{split}&\tilde{\mathbf{y}}:=(0_{D}^{\top},\mathbf{y}_{1:P}^{\top})^{\top},\quad\mathbf{m}=({\Pi_{0}}{\beta}+\Sigma^{-1})^{-1}{\Pi_{0}}{\beta}\tilde{\mathbf{y}}={\beta}({\Pi_{0}}{\beta}+\Sigma^{-1})^{-1}\tilde{\mathbf{y}}\,,\\ &\varphi(\Sigma,\mathbf{y}_{1:P})=\beta{\mathbf{y}_{1:P}^{\top}\mathbf{y}_{1:P}}-(\mathbf{m})^{\top}({\Pi_{0}}{\beta}+\Sigma^{-1})\mathbf{m}\,.\\ \\ \end{split}$
@@
To conclude it remains to compute the previous matrix. We have already noted that, by $\tilde{X}^{\top}\tilde{X}>0$, it follows that also $\Sigma$ is invertible. Being $X^{\top}X$ and $\mathbf{x}_{0}^{\top}\mathbf{x}_{0}$ strictly positive, then the Schur complement of $\Sigma_{00}$ in $\Sigma$, $\Sigma^{*}_{00}:=\Sigma_{00}-\Sigma_{01}\Sigma_{11}^{-1}\Sigma_{01}^{\top}$, is well-defined and strictly positive (Proposition 1.34 Eaton, 2007). In particular, using the Banachiewicz inversion formula, one can write the inverse of $\Sigma$ as
@@
See, e.g., Theorem 1.2 in Zhang (2005). Using again the Banachiewicz inversion formula on $\beta\Pi_{0}+\Sigma^{-1}$ (which is again strictly positive), after some tedious computations, one gets that
@@
$\Sigma_{00}=\frac{\mathbf{x}_{0}^{\top}\mathbf{x}_{0}}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{01}=\frac{\mathbf{x}_{0}^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{11}=\frac{X^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)},$
@@
Proof of Proposition 8. We give the proof in the case $L=1$, the general case follows the same lines, or using Exercise 4.2.7 in Dembo and Zeitouni (2010) and the independence of the components. The result is a consequence of a very general version of the Cramér’s theorem. See Theorem 6.1.3 and Corollary 6.1.6 in Dembo and Zeitouni (2010). Recall that if $Q_{1}\sim\mathcal{W}_{D}(\cdot|\mathbb{1}_{D},N)$ then $Q_{\ell}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}$ where $\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}\sim\mathcal{W}_{D}(\cdot|\mathbb{1}_{D},1)$ are independent and identically distributed taking values in $\mathcal{S}^{+}_{D}$. Now $\mathcal{S}^{+}_{D}$ is a convex open cone in the (topological) vector space $\mathcal{S}_{D}$ of the $D\times D$ symmetric matrices, endowed with the scalar product $(A,B)=\operatorname{tr}(AB)$. The space $\mathcal{S}_{D}$ is locally convex, Hausdorff and the closure of $\mathcal{S}^{+}_{D}$ in $\mathcal{S}_{D}$ is convex and separable and complete (with the topology induced by $\mathcal{S}_{D})$. Clearly $P\{\mathbf{Z}_{j}\mathbf{Z}_{j}^{\top}\in\overline{\mathcal{S}^{+}_{D}}\}=1$ and any closed convex hull of a compact in $\overline{\mathcal{S}^{+}_{D}}$ is compact. Hence all the hypotheses of the General Cramér’s theorem given in Chapter 6 of Dembo and Zeitouni (2010) are satisfied and hence $Q_{\ell}$ satisfies a (weak) LDP (when $N\to+\infty$) with rate function given by $I(Q)=\Lambda^{*}(Q):=\sup_{A\in\mathcal{S}_{D}}\{\operatorname{tr}(QA)-\Lambda(A)\}$ where
@@
In point of fact, see Corollary 6.1.6 in Dembo and Zeitouni (2010), it holds also a strong LDP provided that 00 is in the interior of the domain of $\Lambda$, as we shall show. The well-known expression for the Laplace transform of a Wishart distribution with one degree of freedom (see (20)) gives $\Lambda(A)=-\frac{1}{2}\log(\det(\mathbb{1}_{D}-2A))$ wherever the eigenvalues of $\mathbb{1}_{D}-2A$ are positive and +$\infty$+ ∞ otherwise. In computing $\Lambda^{*}(Q)$ we can restrict $A$ to be such that the eigenvalues of $\mathbb{1}_{D}-2A$ are positive, since otherwise $-\Lambda(A)=-\infty$. Now we check that
@@
Proof. Proposition 10    The proof follows by Lemma  13. To check the assumptions, note that
$Q_{1},\dots,Q_{L}\mapsto\Sigma(Q_{1},\dots,Q_{L})=\begin{pmatrix}\Sigma_{00}&\Sigma_{01}\\ \Sigma_{01}^{\top}&\Sigma_{11}\\ \end{pmatrix}:=\frac{\tilde{X}^{\top}\tilde{X}}{N_{0}\lambda^{*}}\otimes Q^{(L)}$
@@
hence $\Sigma_{11}\mapsto\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$ is bounded on bounded set of $\Sigma_{11}$. Using the continuity of $(Q_{1},\dots,Q_{L})\mapsto\Sigma_{11}(Q_{1},\dots,Q_{L})$ it follows that also $(Q_{1},\dots,Q_{L})\mapsto\log(\det(\mathbb{1}_{DP}+\beta\Sigma_{11}))$ is bounded when $\sum_{\ell=1}^{L}\|Q_{\ell}\|_{2}\leq R$ for any $R$. It remains to check that
@@
Letting $\lambda_{i\ell}$ the eigenvalues of $Q_{\ell}$, the claim follows since $\operatorname{tr}(Q_{\ell})-\log(\det(Q_{\ell}))=\sum_{i=1}^{P}\Big{(}\lambda_{i\ell}+\log(\lambda_{i\ell})\Big{)}$. $\blacklozenge$
@@ A.3 Proofs for the C-DLN
$\begin{split}\bar{\mathbf{s}}_{1:P}^{\top}\mathcal{K}_{\mathrm{C}}(Q_{1},\dots,Q_{L})\bar{\mathbf{s}}_{1:P}&=\sum_{\mu,\nu}\sum_{a_{0}}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},s}}{\lambda^{*}C_{0}}\bar{s}_{\mu}\bar{s}_{\nu}\\ &=\sum_{a_{0}}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{L})_{rs}A^{*}_{r,a_{0}}A^{*}_{s,a_{0}}\end{split}$
@@
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ $=\mathbf{E}\Biggl{[}\det\left((\mathbb{1}_{C_{L}}\otimes K^{(L-1)})^{-1}+\frac{1}{\lambda_{L}C_{L}N_{0}}\mathbb{1}_{C_{L}}\otimes\mathbb{1}_{N_{0}}\otimes(\bar{\mathbf{s}}_{1:P}\bar{\mathbf{s}}_{1:P}^{\top})\right)^{-\frac{1}{2}}$ (32)
$\qquad\qquad\qquad\qquad\det(\mathbb{1}_{C_{L}}\otimes K^{(L-1)})^{-\frac{1}{2}}\Biggr{]}$
@@
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ (33)
$\quad=\mathbf{E}\Biggl{[}\int_{\mathcal{S}^{+}_{N_{0}}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}N_{0}}\sum_{i,j}Q_{L,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}K^{(L-1)}_{ij,\mu\nu}\Biggr{\}}\mathcal{W}_{N_{0}}\Big{(}dQ_{L}|\frac{1}{C_{L}}\mathbb{1}_{N_{0}},C_{L}\Big{)}\Biggr{]}$
$\quad=\mathbf{E}\Biggl{[}\int_{\mathcal{S}^{+}_{N_{0}}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}\lambda_{L-1}C_{L-1}N_{0}}\sum_{i,j}Q^{*}_{L,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}\sum_{a_{L-1}=1}^{C_{L-1}}h^{(L-1)}_{a_{L-1},i,\mu}h^{(L-1)}_{a_{L-1},j,\nu}\Biggr{\}}$
$\qquad\qquad\qquad\qquad\qquad\mathcal{W}_{N_{0}}\Big{(}dQ_{L}|\frac{1}{C_{L}}\mathbb{1}_{N_{0}},C_{L}\Big{)}\Biggr{]}$
@@
$\mathbf{E}[e^{i\bar{\mathbf{s}}_{1:P}^{\top}\mathbf{S}_{1:P}}]$ $=\mathbf{E}\Biggl{[}\int_{(\mathcal{S}^{+}_{N_{0}})^{2}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}\lambda_{L-1}N_{0}}\operatorname{tr}\Biggl{(}Q_{L-1}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}[U^{*}_{L}K^{(L-2)}_{\cdot,\mu\nu}(U^{*}_{L})^{\top}]\Biggr{)}\Biggr{\}}$ (36)
$\qquad\qquad\prod_{\ell\in\{L,L-1\}}\mathcal{W}_{N_{0}}(dQ_{\ell}|\frac{1}{C_{\ell}}\mathbb{1}_{N_{0}},C_{\ell})\Biggr{]}$
$=\mathbf{E}\Biggl{[}\int_{(\mathcal{S}^{+}_{N_{0}})^{2}}\exp\Biggl{\{}-\frac{1}{2\lambda_{L}\lambda_{L-1}\lambda_{L-2}C_{L-2}N_{0}}\sum_{i,j}Q^{*}_{L-1,ij}\sum_{\mu,\nu}\bar{s}_{\mu}\bar{s}_{\nu}$
$\qquad\qquad\sum_{a_{L-2}=1}^{C_{L-2}}h^{(L-2)}_{a_{L-2},i,\mu}h^{(L-2)}_{a_{L-2},j,\nu}\Biggr{\}}\times\prod_{\ell\in\{L,L-1\}}\mathcal{W}_{N_{0}}\Big{(}dQ_{\ell}|\frac{1}{C_{\ell}}\mathbb{1}_{N_{0}},C_{\ell}\Big{)}\Biggr{]}$